OpenStudy (anonymous):

The center of a circle with the equation (x-6)^2 + (y + 2)^2 = 49 is (6, -2). True or False Could you please show me how you got your answer?

jimthompson5910 (jim_thompson5910):

(x-6)^2 + (y + 2)^2 = 49 (x-6)^2 + (y - (-2))^2 = 7^2 notice how it matches the form (x-h)^2+(y-k)^2=r^2 (h,k) is the center so in this problem, h = ?? and k = ??

OpenStudy (anonymous):

@jim_thompson5910 yes, what do H and K come out as?

jimthompson5910 (jim_thompson5910):

well can you see that h = 6 ? if so, then k is what number

OpenStudy (anonymous):

@jim_thompson5910 k=2?

jimthompson5910 (jim_thompson5910):

close, very close

jimthompson5910 (jim_thompson5910):

compare y-k with y - (-2)

OpenStudy (anonymous):

-2?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

h = 6 k = -2

OpenStudy (anonymous):

THANK YOU! You really helped and I learned a lot!

jimthompson5910 (jim_thompson5910):

(h,k) is the center so the center is (6, -2)